# SparsePixels: Efficient Convolution for Sparse Data on FPGAs

### v1

@@ Abstract.
Inference of standard convolutional neural networks (CNNs) on Field-Programmable Gate Arrays (FPGAs) often incurs high latency and a long initiation interval due to the deep nested loops required to densely convolve every input pixel regardless of its feature value, especially when the image size is large. However, in some image data, input features can be spatially sparse, and semantic information may occupy only a small fraction of the input pixels. In this case most computation would be wasted on empty regions. In this work, we introduce SparsePixels, a framework for efficient convolution for spatially sparse image data on FPGAs, targeting fast inference applications in constrained environments with latency requirements of microseconds or below. Our approach implements a special class of CNNs that selectively retain and compute on a small subset of pixels that are active while ignoring the rest. We show that, for example, in a neutrino physics dataset for identifying neutrino interactions in LArTPC images that have around 4k input pixels but are naturally very sparse, a standard CNN with a compact size of 4k parameters incurs an inference latency of 48.665 s on an FPGA, whereas a sparse CNN of the same base architecture computing on less than 1% of the input pixels results in a inference speedup to 0.665 s, with resource utilization well within on-chip budgets, trading only a small percent-level performance loss. At least one-order-of magnitude speedups with comparable performance are also demonstrated in similar datasets with sparse image patterns. This work aims to benefit future algorithm developments for fast and efficient data readout in modern experiments such as the trigger and data acquisition systems at the CERN Large Hadron Collider. For easy adoption, we have developed a library to support building sparse CNNs with quantization-aware training, as well as an HLS implementation for FPGA deployment. Codes are available at https://github.com/hftsoi/sparse-pixels.
@@ 1. Introduction
Proof-of-concept on spatially sparse physics image data, demonstrating significant FPGA inference speedups with comparable model performance.
A Python library for quantization-aware training of sparse CNNs and C++ HLS modules for FPGA deployment.
@@ 5. Summary
We have introduced , a framework for sub-microsecond FPGA inference of CNNs using sparse convolution on spatially sparse 2D data. The key features in our HLS implementation are: (1) dynamically identifying and packing active pixels, a small subset of a large input, into compact arrays, then (2) performing all subsequent computation only on those packed pixels while ignoring inactive pixels. By selectively computing on the necessarily informative subset, the design avoids computation on empty regions, yielding significant inference speedups and resource savings, especially when only percent-level or fewer pixels are active. As a representative example on MicroBooNE neutrino images with 4000 input pixels, a sparse CNN that computes on at most 20 active pixels achieves a inference speedup on an FPGA over a standard CNN with the same base architecture (133 vs. 9733 clock cycles; 0.665 vs. 48.665 s), with ROC AUC drop, and resource utilization well within device budgets (e.g., on the order of 20% LUTs for the largest sparse configuration). For easy adoption, we provide a Python library supporting quantization-aware training of sparse CNNs and HLS C++ modules for FPGA deployment. Codes are available at https://github.com/hftsoi/sparse-pixels, and automated HLS generation will be supported in a future version of . Future directions include implementing sparse layers with stream I/O, exploring partial parallelization, and integrating the framework with high-granularity quantization. We hope this work benefits experiments operating under stringent sub-microsecond latency constraints, such as modern particle experiments, where standard CNNs are often infeasible due to too high a latency and would otherwise force substantial performance compromises.

### v2

@@ Abstract.
Inference of standard convolutional neural networks (CNNs) on FPGAs often incurs high latency and a long initiation interval due to the deep nested loops required to densely convolve every input pixel regardless of its feature value. However, input features can be spatially sparse in some image data, where semantic information may occupy only a small fraction of the pixels and most computation would be wasted on empty regions. In this work, we introduce SparsePixels, a framework that implements sparse convolution on FPGAs by selectively retaining and computing on a small subset of active pixels while ignoring the rest. We show that, for identifying neutrino interactions in naturally sparse LArTPC images with 4k pixels, a standard CNN with a compact size of 4k parameters incurs an inference latency of 48.665 s on an FPGA, whereas a sparse CNN of the same base architecture, computing on less than 1% of the input pixels, achieves a speedup to 0.665 s with resource utilization well within on-chip budgets, trading only a small percent-level performance loss. This work aims to benefit future algorithm development for efficient data readout in modern experiments with latency requirements of microseconds or below. Code is publicly released at https://github.com/hftsoi/sparse-pixels.
@@ 1. Introduction
Proof-of-concept on spatially sparse image data, demonstrating significant FPGA inference speedups with comparable model performance.
A Python library for quantization-aware training of sparse CNNs and C++ HLS modules for FPGA deployment. Code is publicly released at https://github.com/hftsoi/sparse-pixels.
@@ 5. Summary
We have introduced , a framework for sub-microsecond FPGA inference of CNNs using sparse convolution on spatially sparse 2D data. The key features in our HLS implementation are: (1) dynamically identifying and packing active pixels, a small subset of a large input, into compact arrays, then (2) performing all subsequent computation only on those packed pixels while ignoring inactive pixels. By selectively computing on the necessarily informative subset, the design avoids computation on empty regions, yielding significant inference speedups and resource savings, especially when only percent-level or fewer pixels are active. As a representative example on MicroBooNE neutrino images with 4000 input pixels, a sparse CNN that computes on at most 20 active pixels achieves a inference speedup on an FPGA over a standard CNN with the same base architecture (133 vs. 9733 clock cycles; 0.665 vs. 48.665 s), with ROC AUC drop, and resource utilization well within device budgets (e.g., on the order of 20% LUTs for the largest sparse configuration). For easy adoption, we provide a Python library supporting quantization-aware training of sparse CNNs and HLS C++ modules for FPGA deployment. Future directions include implementing sparse layers with stream I/O, exploring partial parallelization, and integrating the framework with high-granularity quantization. We hope this work benefits experiments operating under stringent sub-microsecond latency constraints, such as modern particle experiments, where standard CNNs are often infeasible due to too high a latency and would otherwise force substantial performance compromises.